OpenStudy (anonymous):

-1/8^(2/3) the answer is 1/4 but I would like to know how to solve the problem step by step without using calculators because we are not allowed to use them.

OpenStudy (anonymous):

An expression raised to the a/b power may be rewritten as follows:\[N ^{\frac{ a }{ b }}=\sqrt[b]N ^{a}\] So, in this case your N=-1/8 your b=3 and your a=2. In other words, you are trying to find the cube root of -1/8, and once you find the cube root of -1/8, you take that result and square it.

OpenStudy (anonymous):

Thank you very much I see that you have to take the cube root and then square it to get 1/4

OpenStudy (anonymous):

That is correct, so as long as you have a and b, you can do any problem similar to that.